OpenStudy (anonymous):

Will someone please help me answer this question, What is the discontinuity of the function f(x) = the quantity of x squared minus 3 x minus 28, all over x plus 4

OpenStudy (anonymous):

will you please help me @michelle.k

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

thank you anyway @michelle.k

OpenStudy (anonymous):

someone please help @vshiroky @sarapetrini

OpenStudy (anonymous):

@Rancd please help!!

OpenStudy (anonymous):

@yalo23

OpenStudy (anonymous):

@RebeccaPink6398

OpenStudy (anonymous):

@ninjamicky123

OpenStudy (anonymous):

Read this article before we start: http://www.math.brown.edu/utra/discontinuities.html spik to the part about discontinuity.

OpenStudy (anonymous):

okay now what @Rancd

OpenStudy (anonymous):

So we can write the function like this right ?\[\left( x^{2} - 3\right)\left( - 23 \right)\left( x +4 \right)\]

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

So multiply all this together please

OpenStudy (anonymous):

what do i put in place of x

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

if so i got 276

OpenStudy (anonymous):

No keep x as it is

OpenStudy (anonymous):

i dont know how to do that then

OpenStudy (anonymous):

and multiply if you don't know how to that i will show

OpenStudy (anonymous):

yes please show me

OpenStudy (anonymous):

this will take 3 minutes hold on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

 so we get\[23x^{2} + 69\]